Velocity of a particle is expressed as
Vx=40-5t²m/s
(a)=find average acceleration
(b)=instantaneous acceleration
1
Expert's answer
2020-03-17T09:22:26-0400

(a) The average acceleration


"a_x=\frac{v_2-v_1}{t_2-t_1}=\frac{(40-5t_2^2)-(40-5t_1^2)}{t_2-t_1}=-5(t_2+t_1)."

(b) The instantaneous acceleration


"a_x(t)=v_x'(t)=(40-5t^2)'=-10t."